OpenStudy (anonymous):

solutions of x2 + 2x + 4? (using the quadratic formula!)

OpenStudy (gorv):

(x+2)^2=x^2+4x+4 @dpasingh

OpenStudy (anonymous):

I have to show work using the quadratic formula. I started it but I got lost somewhere.

OpenStudy (anonymous):

Where'd you get lost in the formula

OpenStudy (anonymous):

Ok wait....

OpenStudy (gorv):

USE THIS FORMULA TO SOLVE u will get ur solution

OpenStudy (anonymous):

I did but I messed up somewhere in it...

OpenStudy (anonymous):

Let us compare \[x^2 + 2x + 4=0 \] by \[ax^2 + bx +c=0\] we find: a=1, b=2, c=4 Now let us use the quadratic formula to find its solution. \[x= \frac{-b \pm \sqrt{b^2-4ac}}{2a}= \frac{-2 \pm \sqrt{2^2-4 \times 1\times 4}}{2 \times 1}\] \[= \frac{-2 \pm \sqrt{4-16}}{2}= \frac{-2 \pm \sqrt{-12}}{2}= \frac{-2 \pm 2\sqrt{-3}}{2}\] \[= \frac{2(-1 \pm \sqrt{-3)}}{2}=-1 \pm \sqrt{-3} \] is the required solution of the given quadratic equation. @brooklynransome

OpenStudy (gorv):

@brooklynransome now he is absolutely right

OpenStudy (anonymous):

@brooklynransome Your problem has been solved.